# Vulnerability of Text-Matching in ML/AI Conference Reviewer Assignments to Collusions

### v1

@@ 1 Introduction
The rapid growth of machine learning and artificial intelligence research has led to major scientific conferences in these fields receiving thousands to tens of thousands of paper submissions and similar numbers of reviewers [Sha22]. Consequently, the number of reviewers is also in the thousands or tens of thousands. To manage this overwhelming volume, the assignment of reviewers to papers has become largely automated. A key component of this automated assignment pertains to text-matched similarity scores between reviewers’ past work and submitted manuscripts, where natural language processing algorithms compute the similarity between the text of each submitted manuscript and the texts of the reviewer’s previously published manuscripts. Models like SPECTER [CFB+20], which generates embeddings based on textual content, have become widely used in this automated reviewer assignment process across various prestigious venues, including NeurIPS, and are a default model on the popular OpenReview conference management platform.
@@
Although automation facilitates efficient handling of large-scale peer review, it also introduces potential vulnerabilities. One major challenge threatening the integrity of the peer review process is the existence of collusion rings – groups of individuals who conspire to manipulate the review system for personal gain [Lit21, Vij20a]. These rings can unfairly influence the acceptance of certain papers by orchestrating favorable reviews from colluding reviewers.
A critical tactic employed by collusion rings is manipulating the reviewer assignment process to ensure that allied reviewers are assigned to specific colluding papers. For example, as noted in [Vij20a], colluders may attempt to “game the system by exploiting vulnerabilities in the assignment algorithms to have their collaborators review their submissions.” Research and practical efforts have thus far focused on detecting and mitigating collusion through the analysis of bidding, the part of the reviewer assignment process where reviewers can indicate their interest in reviewing each paper.
@@
Studies such as those by [WGW+21, JSFA24] to identify or mitigate suspicious activities focus on bidding patterns. The algorithm in [WGW+21] considers text similarities as ground truth when examining the issues related to bids. In practice, program chairs have also extensively focused on bidding patterns when investigating collusion rings [PC24]; in response to concerns over bidding manipulation, venues such as CVPR and ARR have removed the bidding process altogether. This focus on bidding implicitly or explicitly assumes that the text-matching algorithms used for reviewer assignments are secure and resistant to manipulation. However, this assumption warrants scrutiny.
@@
In this paper, we investigate whether the text-matching algorithms used in automated reviewer assignments are indeed robust against manipulation. Since most papers are assigned to 3–6 reviewers at conferences [Sha22], we focus our evaluation on whether a pair of colluding author and reviewer can successfully manipulate the text-matching algorithms to give the colluding reviewer one of the top-1, top-3, or top-5 highest text similarity to the colluding paper among all reviewers at a conference. Our findings reveal that these text matching algorithms used in automated reviewer assignments are susceptible to attacks that can significantly increase the calculated similarity between a paper and its colluding reviewer. Here are some salient details:
@@
  • We find that the SPECTER [CFB+20] text similarity matching algorithm, used by various top ML/AI conferences for their reviewer assignments, can be manipulated by colluding authors and reviewers. In evaluations on NeurIPS 2023 data, the human-in-the-loop version of our attack successfully increases a colluding reviewer’s similarity from ranked 101 to top-5 for a colluding paper 92% of the time; in the fully automatic, our attack successfully increases a colluding reviewer’s similarity from ranked 101 to top-5 for a colluding paper 88% of the time.
@@
  • Considering NeurIPS 2022 as a “past” conference whose data is publicly available to an attacker, and NeurIPS 2023 as the “current” target conference whose data is unavailable, we find that the performance of the attack on NeurIPS 2022 is reflective of its performance in NeurIPS 2023. We find strong correlation ($\rho\geq 0.85$) between colluding reviewers’ similarity rankings in NeurIPS 2022 and NeurIPS 2023.
@@
  • Many conferences allow reviewers to select a subset of papers in their profile. We find that a careful selection of the past papers by the reviewer – and particularly choice of fewer papers – can significantly increase the attack’s success rate. In particular, when a colluding reviewer selects only one paper that is the most similar to the colluding paper, it increases the colluding reviewer’s similarity from ranked 101 to top-5 for the colluding paper 42% of the time even without any modifications to the colluding paper’s abstract.
@@
We release our code, datasets, and manipulation examples at:
https://github.com/passionfruit03/reviewer_assignments_vulnerability.
This paper aims to make the community aware of this vulnerability, which has implications for the fairness and integrity of the peer review process. Understanding and addressing these vulnerabilities is crucial to maintaining trust in peer review. In the final section of the paper, we discuss the implications of our findings and propose recommendations to enhance the security and robustness of automated reviewer assignment systems.
@@ 2 Problem setting
In this section, we describe the background on automated reviewer assignments that are widely used in ML/AI conferences, along with our problem setting including the attacker’s threat model.
@@ 2.1 Automated Reviewer Assignments
We consider peer review at conferences, which are the primary venue of publication in computer science. In conference review, there is a pre-chosen pool of reviewers, and we denote this set of all reviewers as $\mathcal{R}$. We let $\mathcal{P}$ denote the set of all submitted papers. In order to assign reviewers to papers, the program chairs compute a “similarity score” for each reviewer-paper pair. The similarity score is a number between 0 and 1, and a higher value of the similarity score indicates a higher envisaged expertise of that reviewer for that paper.
The similarity score is often computed as a combination of multiple distinct components. A primary component is text matching between the submitted paper and the reviewer’s previous papers, which is the main focus of this paper and is discussed in more detail below. Another important component is bidding, where each reviewer can see the list of submitted papers and indicate their interest in reviewing each paper. Prior studies [JZL+20, WGW+21] have demonstrated that bidding is easily susceptible to manipulation. Most previous research on guarding against collusion rings [JZL+20, WGW+21, JSFC22, JYC+23, JSFA24, LBNZ+24], as well as practical implementations (e.g., [PC24, LBNZ+24]), have focused on bidding. For example, peer-reviewed venues like the Conference on Computer Vision and Pattern Recognition (CVPR) and ACL Rolling Review (ARR) have completely disabled bidding to prevent manipulation. Furthermore, it is generally assumed that bidding can be gamed, while text similarity-based methods are considered robust (e.g., in [WGW+21], as well as by venues that have banned bidding in favor of text similarity approaches). A third component involves the subject areas or keywords provided by the authors and reviewers, though it is known that these can also be gamed [ACDK19].
Given the similarity scores, reviewers are then assigned to papers by solving an optimization problem aimed at choosing the assignment with the highest similarity scores, subject to various constraints, such as reviewer and paper loads and known conflicts of interest. Here, a higher similarity score for any reviewer-paper pair means a greater chance of this reviewer being assigned to this paper. Different conferences employ different optimization programs [CZ13b, SSS21, KSM19, JZL+20, PZ22, LBNZ+24]. To understand whether reviewer assignments can be manipulated for a colluding paper-reviewer pair, we evaluate the increase in the colluding reviewer’s similarity score ranking when compared to other reviewers for the paper, so the results of this paper are generally applicable to different optimization problems specifically defined by each venue.
With this context, we focus on the vulnerability of the similarity score that depends only on text similarity, since bidding is known to be vulnerable and can only increase the success of reviewer assignment manipulations.
@@ 2.2 Paper-Reviewer Text Similarity
For a $p\in\mathcal{P}$ and reviewer $r\in\mathcal{R}$, we let $s({p},{r})\in[0,1]$ denote the computed text similarity between this reviewer and this paper. This similarity can be computed in many ways [CZ13b, CFB+20, HSC21, SDC+22, ORA+22]. In this work, we focus on the widely used SPECTER algorithm [CFB+20].
SPECTER is widely used for assigning reviewers in ML/AI related venues, for instance, in conferences such as NeurIPS, ICML, ICLR and the TMLR journal. SPECTER leverages neural network embeddings that are trained with an objective to keep the embeddings of similar papers close, where two papers are considered similar when one of them cites the other.
For any reviewer $r\in\mathcal{R}$, we let $\mathcal{Q}_{r}$ denote as the archive containing $10$ most recent papers authored by $r$. The similarity score $s({p},{r})$ is an aggregation of the cosine similarities between the SPECTER embedding of $p$ and the SPECTER embeddings for each of the different papers in $\mathcal{Q}_{r}$. Formally, for any paper $p$, let $\mathbf{v}_{p}\in\mathbb{R}^{768}$ denote its SPECTER embedding vector. In practice, only the titles and abstracts are used to compute $\mathbf{v}_{p}$. Then, for a submitted paper $p\in\mathcal{P}$ and each paper $q\in\mathcal{Q}_{r}$ in a reviewer’s archive, the individual similarity between $p$ and $q$ is computed as the cosine similarity between their embeddings: $\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$.
The protocol for handling multiple papers in the reviewer’s archive $\mathcal{Q}_{r}$ involves the following two steps: (i) A similarity score is computed between the submitted paper and each individual paper in the reviewer’s archive (described above); and (ii) To determine an overall similarity score between the submitted paper and the reviewer, the individual similarity scores across the papers in the reviewer’s archive are then aggregated. Two commonly used methods for this aggregation are:
  • Average (or mean) pooling: $s({p},{r})=\frac{1}{\lvert\mathcal{Q}_{r}\rvert}\underset{q\in\mathcal{Q}_{r}}{\sum}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$, and
  • Max pooling: $s({p},{r})=\underset{q\in\mathcal{Q}_{r}}{\max}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$.
Max pooling and average pooling are both commonly used, with max pooling generally preferred over average pooling because it has been observed to provide more accurate similarity scores.
In addition, some conferences may have $\mathcal{Q}_{r}$ containing a different number of most recent papers by default. In fact, it is common to allow reviewers to curate their own archives, which as we demonstrate provides an attack surface.
@@ 2.3 Attacker’s Threat Model
The “attacker” represents both the colluding author of a paper $p\in\mathcal{P}$ and a reviewer $r\in\mathcal{R}$ working together to ensure $p$ is assigned to $r$ for peer review. For text-similarity-based assignments that we consider in this work, this corresponds to increasing $s({p},{r})$ so that $r$ ranks in top-k of the conference’s reviewers ranked by their similarity scores with $p$. In our subsequent experiments, we consider $k\in\{1,3,5\}$.
@@
We assume that each pair of colluding author and reviewer (the “attacker”) know each other and are actively colluding, so they can work together to create the attack. The colluding author can manipulate the abstract of their paper, and the colluding reviewer can adversarially curate their reviewer archive $\mathcal{Q}_{r}^{\text{adv}}$ to only retain selected papers. We only consider abstract modifications, since the SPECTER similarity scores are computed based on title and abstract only, though the author can change any parts of their paper. In this paper, we also do not consider changes to the title because we suspect title modifications may arouse more suspicion from non-colluding reviewers.
When modifying the abstract, the attacker must avoid arousing suspicion in non-colluding reviewers. This is ostensibly vague and we formalize this notion via two constraints on adversarial modifications of the abstract: (i) Coherence: The abstract should use academic writing style, have natural flow, and cannot contain scientifically false information. (ii) Consistency: The abstract should be consistent with the paper’s contents. In this paper, we also enforce constraints on the number of sentences and keywords that can be added to the colluding paper’s abstract. Attacks are generally allowed only one sentence that is related to $\mathcal{Q}_{r}^{\text{adv}}$, but we allow up to three sentences related to $\mathcal{Q}_{r}^{\text{adv}}$ in some cases only when there is manual supervision involved to ensure coherence and consistency. In addition, attacks are allowed to add up to 10 keywords selected from $\mathcal{Q}_{r}^{\text{adv}}$. We also perform human-study experiments to judge whether non-colluding reviewers find that abstracts generated by our proposed attack are suspicious.
@@
The attackers have access to the exact SPECTER embeddings that are publicly available at https://github.com/allenai/specter. For many conference venues, the attackers also have access to the publicly available reviewer pool from the previous year’s conference (e.g., https://neurips.cc/Conferences/2023/ProgramCommittee). However, they do not have access to the entire reviewer pool of the current iteration of the conference.
@@ 3 Related work
Our focus is the text matching between reviewers and papers. There are a number of algorithms for computing text matching similarity scores [MM07, CZ13a, WGNBK19, CFB+20, MJMZ23, SDC+22]; see [Sha22, Section 3] for a more extensive survey. In this work we focus on the SPECTER model [CFB+20], described in Section 2.2, due to its widespread use. We note that a second version of SPECTER is also available [SDC+22]. In preliminary experiments, we found that SPECTER version 2 has similar vulnerabilities as the original SPECTER model, and we focus on SPECTER since this is still the most widely used model in ML/AI conferences.
@@
    Setting While the demonstration of manipulation in their work is on a small scale of 165 reviewers and 32 papers, our work considers a scale two orders of magnitude larger with 7,900 reviewers and 3,218 papers. We also consider attacks when the original ranking of the reviewer for the target paper is in the range of 101 to 1001, as compared to the original ranking of 6 to 10 considered in their work. An additional assumption made in their work is that the attackers know the program committee beforehand. This assumption is not suitable for our setting (ML/AI venues), so we also dispose of this assumption. Finally, their work focuses only on modifying the paper submission, but our work additionally considers adversarial archive curation by the reviewer as another attack surface.
@@
    Text matching model Their work considers a Latent Dirichlet Allocation (LDA) based topic modeling algorithm used in the Toronto Paper Matching System (TPMS) for text matching. In our work, we instead consider a modern neural network embedding based text similarity model [CFB+20]. The model we consider is much less interpretable than LDA-based models, and it is more widely used, deployed in most top-tier ML/AI venues. The paper [EQM+23] further investigates the possibility of black box attacks without access to the reviewer assignment model. However, in our scenario, we assume that anyone has access to the open-sourced SPECTER model.
@@
    Evaluation The manipulations in their work involve various PDF tricks, without which they achieve relatively low success rates ($<70\%$ even with aggressive modifications to the paper submission); in contrast, we show that assignment manipulation purely through text manipulations is possible with high success rates ($\sim 88\%$). In our work, we focus on investigating the possibility of a paper submission getting assigned to a colluding reviewer, without considering adding or removing multiple reviewers at once. In their ablation, they demonstrated that it is possible to select all five reviewers adversarially using PDF manipulations to add or remove a median of 5,968 words from the submissions. Finally, while both papers include human subject experiments to test for the detectability of adversarial papers, we collected 116 samples, which is considerably more than the 21 samples they had collected.

### v2

@@ 1 Introduction
The rapid growth of machine learning and artificial intelligence research has led to major scientific conferences in these fields receiving thousands to tens of thousands of paper submissions and similar numbers of reviewers [Sha22]. To manage this overwhelming volume, the assignment of reviewers to papers has become largely automated. A key component of this automated assignment pertains to text-matched similarity scores between reviewers’ past work and submitted manuscripts, where natural language processing algorithms compute the similarity between the text of each submitted manuscript and the texts of the reviewer’s previously published manuscripts. Models like SPECTER [CFB+20], which generates embeddings based on textual content, have become widely used across various prestigious venues, including the Conference on Neural Information Processing Systems (NeurIPS), and are a default model on the popular OpenReview conference management platform. [^1] While our work focuses on the peer-review process in ML/AI venues, the continuous evolution and scaling up of other communities suggest that they may soon have to adopt similar processes and face similar challenges.
@@
Although automation facilitates efficiency, it also introduces potential vulnerabilities. One major challenge threatening the integrity of the peer review process is the existence of collusion rings – groups of individuals who conspire to manipulate the review system for personal gain [Lit21, Vij20a]. These rings can unfairly influence the acceptance of certain papers by orchestrating favorable reviews from colluding reviewers. As noted in [Vij20a], colluders may attempt to “game the system by exploiting vulnerabilities in the assignment algorithms to have their collaborators review their submissions.”
@@
Studies to identify or mitigate collusion have thus far focused on analyzing manipulations in bidding, the part of the reviewer assignment process where reviewers can indicate their interest in reviewing each paper [WGW+21, JSFA24]. The algorithm in [WGW+21] considers text similarities as ground truth when examining the bids. In practice, program chairs have also focused on bidding when investigating collusion rings [PC24]; to address concerns over bidding manipulation, venues such as the Conference on Computer Vision and Pattern Recognition (CVPR) and ACL Rolling Review (ARR) have removed the bidding process altogether. This focus on bidding implicitly or explicitly assumes that the text-matching algorithms resistant to manipulation. However, this assumption warrants scrutiny.
@@
In this paper, we investigate whether the text-matching algorithms used in automated reviewer assignments are indeed robust against manipulation. Since most papers are assigned to 3–6 reviewers at conferences [Sha22], we focus our evaluation on whether a pair of colluding author and reviewer can successfully manipulate the text-matching algorithms to give the colluding reviewer one of the top-1, top-3, or top-5 highest text similarity to the colluding paper among all reviewers at a conference. We find that the text matching algorithms are susceptible to attacks that can significantly increase the calculated similarity between a paper and its colluding reviewer. Here are some salient details:
@@
  • We find that the SPECTER [CFB+20] text similarity matching algorithm, used by top ML/AI conferences for reviewer assignments, can be manipulated by colluding authors and reviewers. In evaluations on NeurIPS 2023 data, our attack successfully increases a colluding reviewer’s similarity from ranked 101 to top-5 for a colluding paper about $92\%$ of the time.
@@
  • Considering NeurIPS 2022 as a “past” conference whose data is publicly available to an attacker, and NeurIPS 2023 as the “current” conference whose data is unavailable, we find that the attack performance on NeurIPS 2022 is reflective of the performance in NeurIPS 2023. We find strong correlations of 0.62, 0.83, 0.92 and 0.93 between colluding reviewers’ similarity rankings in NeurIPS 2022 and NeurIPS 2023 in four different settings.
@@
  • Many conferences allow reviewers to select a subset of papers in their profile. We find that a careful selection of the past papers by the reviewer – and particularly a choice of fewer papers – can significantly increase the attack’s success rate. When a colluding reviewer selects only one paper that is the most similar to the colluding paper, it increases the colluding reviewer’s similarity from ranked 101 to top-5 for the colluding paper 41% of the time even without any modifications to the colluding paper’s abstract.
@@
In the final section of the paper, we discuss the implications of our findings and propose recommendations to enhance the security and robustness of automated reviewer assignments systems. We have also informed the concerned conference management platform about these results. Based on the findings of this work, the platform has implemented safeguards for conference organizers to guard against such attacks, and the safeguards have been used by a major AI/ML conference. Subsequent peer-review venues have also expressed interest in deploying safeguards, and we are closely helping these venues in doing so.
We release our research artifacts including code, datasets, and manipulation examples on Zenodo https://doi.org/10.5281/zenodo.15588237. This paper aims to make the community aware of this vulnerability, which has implications for the fairness and integrity of the peer review process.
@@ 2 Problem setting
In this section, we describe the automated reviewer assignments that are widely used in ML/AI conferences, along with our problem setting and the attacker’s threat model.
@@ 2.1 Automated Reviewer Assignments
We consider peer review at conferences, which are the primary venue of publication in computer science. In conference review, there is a pre-chosen pool of reviewers, and we denote this set of all reviewers as $\mathcal{R}$. We let $\mathcal{P}$ denote the set of all submitted papers. To assign reviewers to papers, the program chairs compute a “similarity score” for each reviewer-paper pair. The similarity score is a number between 0 and 1, and a higher value of the similarity score indicates a higher envisaged expertise of that reviewer for that paper.
The similarity score is often a combination of multiple components. A primary component is text matching between the submitted paper and the reviewer’s previous papers, which is the focus of this paper and is discussed below. Another important component is bidding, where each reviewer can see the list of submitted papers and indicate their interest in reviewing each paper. Prior studies [JZL+20, WGW+21] have demonstrated that bidding is easily susceptible to manipulation. Most previous research on guarding against collusion rings [JZL+20, WGW+21, JSFC22, JYC+23, JSFA24, LBNZ+24], as well as practical implementations (e.g., [PC24, LBNZ+24]), have focused on bidding. For example, venues like CVPR and ARR have completely disabled bidding to prevent manipulation. Furthermore, it is generally assumed that bidding can be gamed, while text similarity-based methods are considered robust (e.g., in [WGW+21], as well as by venues that have banned bidding in favor of text similarity approaches). A third component involves the subject areas or keywords provided by the authors and reviewers, though these can also be gamed [ACDK19].
Reviewers are then assigned to papers by solving an optimization problem to choose the assignment with the highest similarity scores, subject to various constraints, such as reviewer and paper loads and known conflicts of interest. A higher similarity score for any reviewer-paper pair means this reviewer is more likely to get assigned the paper. To answer if reviewer assignments can be manipulated for a colluding paper-reviewer pair, we evaluate the increase in the colluding reviewer’s similarity score ranking when compared to other reviewers for the paper. Since different conferences have different optimization programs [CZ13b, SSS21, KSM19, JZL+20, PZ22, LBNZ+24], evaluating the similarity rankings, which are relative, makes our results more generally applicable. With this context, we focus on the vulnerability of the similarity score that depends only on text similarity, since bidding is known to be vulnerable and can only increase the success of reviewer assignment manipulations.
@@ 2.2 Paper-Reviewer Text Similarity
For a $p\in\mathcal{P}$ and reviewer $r\in\mathcal{R}$, we let $s({p},{r})\in[0,1]$ denote the text similarity between this reviewer and paper. This similarity can be computed in many ways [CZ13b, CFB+20, HSC21, SDC+22, ORA+22]. In this work, we focus on the widely used SPECTER algorithm [CFB+20].
SPECTER is widely used for assigning reviewers in ML/AI related venues such as NeurIPS, the International Conference on Machine Learning (ICML), the International Conference on Learning Representations (ICLR) and the Transactions on Machine Learning Research (TMLR) journal. SPECTER uses neural network embeddings trained with an objective to keep the embeddings of similar papers close, where two papers are considered similar when one of them cites the other.
For any reviewer $r\in\mathcal{R}$, we let $Q_{r}$ denote as the default archive containing $10$ most recent papers authored by $r$. The similarity score s(p,r)s({p},{r}) is an aggregation of the cosine similarities between the SPECTER embedding of $p$ and the SPECTER embeddings for each of the different papers in $Q_{r}$. Formally, for any paper $p$, let $\mathbf{v}_{p}\in\mathbb{R}^{768}$ denote its SPECTER embedding vector. In practice, only the titles and abstracts are used to compute $\mathbf{v}_{p}$. Then, for a submitted paper $p\in\mathcal{P}$ and each paper $q\in Q_{r}$ in a reviewer’s archive, the individual similarity between $p$ and $q$ is computed as the cosine similarity between their embeddings: $\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$.
The protocol for handling multiple papers in the reviewer’s archive $Q_{r}$ involves the following two steps: (i) A similarity score is computed between the submitted paper and each individual paper in the reviewer’s archive (described above); and (ii) To determine an overall similarity score between the submitted paper and the reviewer, the individual similarity scores across the papers in the reviewer’s archive are then aggregated. Two commonly used aggregation methods are:
  • Average (mean) pooling: $s({p},{r})=\frac{1}{\lvert Q_{r}\rvert}\underset{q\in Q_{r}}{\sum}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$, and
  • Max pooling: $s({p},{r})=\underset{q\in Q_{r}}{\max}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}$.
Max pooling and average pooling are both commonly used, with max pooling generally preferred because it has been observed to provide more accurate similarity scores.
In addition, some conferences may have $Q_{r}$ containing a different number of most recent papers by default. In fact, it is common to allow reviewers to curate their own archives, which as we demonstrate provides an attack surface.
@@ 2.3 Attacker’s Threat Model
The “attacker” represents both the colluding author of a paper $p\in\mathcal{P}$ and a reviewer $r\in\mathcal{R}$ working together to ensure $p$ is assigned to $r$ for peer review. For text-similarity-based assignments, this corresponds to increasing s(p,r)s({p},{r}) so that $r$ ranks in top-k of the conference’s reviewers ranked by their similarity scores with $p$. In our subsequent experiments, we consider $k\in\{1,3,5\}$.
@@
We assume that each pair of colluding author and reviewer (the “attacker”) know each other and are actively colluding, so they can work together to create the attack. The colluding author can manipulate the abstract of their paper, and the colluding reviewer can adversarially curate their reviewer archive $Q_{r}^{\text{adv}}$ to only retain selected papers. We only consider abstract modifications, since the SPECTER similarity scores are computed based on title and abstract only, though the author can change any parts of their paper. In this paper, we also do not consider changes to the title because we suspect title modifications may arouse more suspicion from non-colluding reviewers.
When modifying the abstract, the attacker must avoid arousing suspicion in non-colluding reviewers. This is ostensibly vague and we formalize this notion via two constraints on adversarial modifications of the abstract: (i) Coherence: The abstract should use academic writing style, have natural flow, and cannot contain scientifically false information. (ii) Consistency: The abstract should be consistent with the paper’s contents. In this paper, we also enforce constraints on the number of sentences and keywords that can be added to the colluding paper’s abstract. Attacks are generally allowed only one sentence that is related to $Q_{r}^{\text{adv}}$, but we allow up to three sentences related to $Q_{r}^{\text{adv}}$ in some cases only when there is manual supervision involved to ensure coherence and consistency. In addition, attacks are allowed to add up to 10 keywords selected from $Q_{r}^{\text{adv}}$. We also perform human-study experiments to judge whether non-colluding reviewers find that abstracts generated by our proposed attack are suspicious.
@@
The attackers have access to the exact SPECTER embeddings, with the model weights publicly available at https://github.com/allenai/specter. For many conference venues, the attackers also have access to the publicly available reviewer pool from the previous year’s conference (e.g., https://neurips.cc/Conferences/2023/ProgramCommittee). However, they do not have access to the reviewer pool of the current iteration of the conference.
@@ 3 Related work
There are a number of algorithms for computing text matching similarity scores [MM07, CZ13a, WGNBK19, CFB+20, MJMZ23, SDC+22]; see [Sha22, Section 3] for a more extensive survey. In this work we focus on the SPECTER model [CFB+20], described in Section 2.2, due to its widespread use. We note that a second version of SPECTER is also available [SDC+22]. In preliminary experiments, we found that SPECTER version 2 has similar vulnerabilities as the original SPECTER model, and we focus on SPECTER since this is still the most widely used model in ML/AI conferences.
@@
    Setting While the demonstration of manipulation in their work is on a small scale of 165 reviewers and 32 papers, our work considers a scale two orders of magnitude larger with 7,900 reviewers and 3,218 papers. We also consider attacks when the original ranking of the reviewer for the target paper is in the range of 20 to 1001, as compared to the original ranking of 6 to 10 considered in their work. An additional assumption made in their work is that the attackers know the program committee beforehand. This assumption is not suitable for our setting (ML/AI venues), so we also dispose of this assumption. Finally, their work focuses only on modifying the paper submission, but our work additionally considers adversarial archive curation by the reviewer as another attack surface.
@@
    Text matching model Their work considers a Latent Dirichlet Allocation (LDA) based topic modeling algorithm used in the Toronto Paper Matching System (TPMS) for text matching. In our work, we instead consider a modern neural network embedding based text similarity model [CFB+20]. The model we consider is much less interpretable than LDA-based models, and it is more widely used at ML/AI venues. The paper [EQM+23] further investigates the possibility of black box attacks without access to the reviewer assignment model. However, in our scenario, we assume that anyone has access to the open-sourced SPECTER model.
@@
    Evaluation The manipulations in their work involve PDF tricks, without which they achieve relatively low success rates ($70\%$ even with aggressive modifications to the paper); in contrast, we show that assignment manipulation purely through text manipulations is possible with high success rates ($\sim 92\%$). In our work, we focus on investigating the possibility of a paper submission getting assigned to a colluding reviewer, without considering adding or removing multiple reviewers at once. In their ablation, they demonstrated that it is possible to select all five reviewers adversarially using PDF manipulations to add or remove a median of 5,968 words from the submissions. Finally, while both papers include human subject experiments to test for the detectability of adversarial papers, we collected 116 samples, which is considerably more than the 21 samples they collected.